This question may apply more to a pre-calc class, but this may be the closes subject.
simplify the expression:
(cos((pi/2)-x)tanx)/(sin^2x)
Please Help!!!

Found 2 solutions by richard1234, lwsshak3:
Answer by richard1234(7193)



Since cos (pi/2 - x) = sin x, the expression is equal to



By definition, tan x = sin x/cos x, replace tan x.


Answer by lwsshak3(11628)

simplify the expression:
(cos((pi/2)-x)tanx)/(sin^2x)
cos(π/2-x)=sinxx
sin*(sin/cos)/sin^2
=sin^2/cos*sin^2
=1/cos=sec